# Solving Inverse Problems with FLAIR

### v1

@@ 5.3 Ablation Studies
We systematically analyze the impact of key design choices in our method. Specifically, we ablate the deterministic trajectory adjustment, the use of hard data consistency, and the calibration of the regularizer weight for ×$\times 12$ super-resolution, using a subset of 100 samples from the FFHQ and DIV2K datasets. Quantitative and qualitative results are shown in Table 2 and Figure 5, respectively.
Hard Data Consistency (HDC). Dropping the hard data consistency degrades both metrics, with PSNR being particularly affected due to poorer alignment with the input observation, which is also evident in the visual example: the reconstruction is plausible, but deviates from the observation.
Deterministic Trajectory Adjustment (DTA). The biggest performance drop compared to the full setup occurs when removing the deterministic trajectory adjustment, as standard noise injection harms the gradient updates in low-density regions of the prior. The reconstruction appears overly smooth and lacks texture details.
@@ 6 Conclusion and Limitations
We have presented FLAIR, a training-free variational framework for inverse problems that uses a flow-based generative model as its image prior. By combining the power of (latent) flow-based models with a principled reconstruction of the posterior distribution, FLAIR addresses key limitations of existing methods. First, it is able to recover atypical modes, by combining the degradation-agnostic flow matching loss with deterministic trajectory adjustments. Second, it enables hard data consistency without sacrificing sample diversity, by decoupling the data consistency constraint from the regularization, while adaptively reweighting the latter according to its expected accuracy, calibrated offline. Experiments with different image datasets and tasks confirm that FLAIR consistently achieves higher reconstruction quality than existing baselines based on either flow matching or denoising diffusion. Notably, our proposed method achieves, at the same time, excellent perceptual quality, close adherence to the input observations, and high sample diversity.
Evidently, FLAIR inherits the limitations of the underlying generative model, in our implementation Stable Diffusion 3. These include biases caused by the selection of training data, constraints w.r.t. the output resolution, and a limited ability to recover out-of-distribution modes. Furthermore, our approach introduces additional hyper-parameters needed to control the deterministic trajectory adjustment. We note that high fidelity image restoration methods can potentially be misused for unethical image manipulations.
@@ A.1 Derivation of flow-based variational formulation
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{q(x_{t}|y)}\left[v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right]\;.$ (25)
@@ A.3 Derivation of Score from Flow
[…]
@@
$-\frac{1}{\sigma(t)^{2}}\left(x_{t}-\mu(x_{0},t)\right)=\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{0}),$ (35)
@@ A.5 Baselines
FlowChef [35] Additionally, [24] is employed for FlowChef as well, using 50 NFE, a classifier-free guidance scale of 2, and a step size of 1 for all tasks.
@@ A.8 Additional Experimental Results
Table 5: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$ – Super-resolution (×$\times 8$ and ×$\times 12$).
SR ×$\times 8$ SR ×$\times 12$
@@
A.9 Additional Qualitative Examples
@@
Figure 8: ×$\times 12$ super-resolution results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
@@
Figure 11: ×$\times 12$ super-resolution results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.

### v2

@@ 5.2 Experimental Results
[…]
Figure 5: Qualitative comparison. FLAIR in pixel space produces posterior samples of high perceptual quality while maintaining high data likelihood as well. Best viewed zoomed in.
Table 2: Quantitative results with 50 NFE and $\sigma_{\nu}=0.5\%$ – In-painting and Super-resolution ($\times 8$).
Inpainting SR $\times 8$
[…]
DDNM 0.158 $\pm$ 0.042 26.9 0.732 $\pm$ 0.037 18.31 $\pm$ 2.94 0.199 $\pm$ 0.052 31.9 0.635 $\pm$ 0.079 23.59 $\pm$ 1.64
DPS 0.195 $\pm$ 0.064 30.2 0.689 $\pm$ 0.077 20.49 $\pm$ 2.81 0.172 $\pm$ 0.058 27.8 0.658 $\pm$ 0.088 24.59 $\pm$ 2.04
$M$ 0.161 $\pm$ 0.054 28.8 0.728 $\pm$ 0.062 20.59 $\pm$ 3.37 0.172 $\pm$ 0.051 29.1 0.669 $\pm$ 0.083 24.65 $\pm$ 1.97
[…]
FLAIR 0.097 $\pm$ 0.035 14.2 0.831 $\pm$ 0.031 21.87 $\pm$ 2.66 0.143 $\pm$ 0.039 22.9 0.712 $\pm$ 0.076 25.93 $\pm$ 1.96
@@ 5.3 Ablation Studies
We systematically analyze the impact of key design choices in our method. Specifically, we ablate the deterministic trajectory adjustment, the use of hard data consistency, and the calibration of the regularizer weight for $\times 12$ super-resolution, using a subset of 100 samples from the FFHQ and DIV2K datasets. Quantitative and qualitative results are shown in Table 3 and Figure 6, respectively.
Hard Data Consistency (HDC). Dropping the hard data consistency degrades both metrics, with PSNR being particularly affected due to poorer alignment with the input observation, which is also evident in the visual example: the reconstruction is plausible but deviates from the observation.
Deterministic Trajectory Adjustment (DTA). The biggest performance drop compared to the full setup occurs when removing the deterministic trajectory adjustment, as random noise sampling harms the gradient updates in low-density regions of the prior. The reconstruction appears overly smooth and lacks texture details.
@@
Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗Refer to captionGround truthRefer to captionHDC✓, DTA✓, CRW✓Refer to captionHDC✗, DTA✓, CRW✓Refer to captionHDC✓, DTA✗, CRW✓Refer to captionHDC✓, DTA✓, CRW✗Refer to captionHDC✗, DTA✗, CRW✗
@@ 6 Conclusion and Limitations
We have presented FLAIR, a training-free variational framework for inverse problems that uses a flow-based generative model as its image prior. By combining the power of (latent) flow-based models with a principled reconstruction of the posterior distribution, FLAIR addresses key limitations of existing methods. First, it is able to target the generation towards images, which match the observation, by aiding the degradation-agnostic flow matching loss with deterministic noise vectors. Second, it enables hard data consistency without sacrificing sample diversity, by decoupling the data consistency constraint from the regularization, while adaptively reweighting the latter according to its expected accuracy, calibrated offline. Experiments with different image datasets and tasks confirm that FLAIR consistently achieves higher reconstruction quality than existing baselines based on either flow matching or denoising diffusion. Notably, our proposed method achieves, at the same time, excellent perceptual quality, close adherence to the input observations, and high sample diversity.
Evidently, FLAIR inherits the limitations of the underlying generative model. These include biases caused by the selection of training data, constraints w.r.t. the output resolution, and a limited ability to recover out-of-distribution modes. Furthermore, our approach introduces additional hyper-parameters needed to control the deterministic trajectory adjustment. We note that high fidelity image restoration methods can potentially be misused for unethical image manipulations.
Acknowledgments This work was funded, in part, by Huawei Technologies Oy (Finland) Co. Ltd.
@@ A.1 Derivation of flow-based variational formulation
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{t,q(x_{t}|y)}\left[v_{\theta}(x_{t},t)-u_{t}(x_{t}|\epsilon)\right]\;.$ (26)
@@ A.3 Derivation of Score from Flow
$\nabla_{x_{t}}\ln p_{t}(x_{t})=\underset{\theta}{\arg\min}~\mathbb{E}_{t\sim\mathcal{U}[0,1],x_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,I)}\left[w(t)\cdot\left\|s_{\theta}(x_{t},t)+\frac{1}{\sigma(t)^{2}}\left(x_{t}-\mu(x_{0},t)\right)\right\|^{2}\right],$ (35)
@@
$\frac{1}{\sigma(t)^{2}}\left(x_{t}-\mu(x_{0},t)\right)=\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{0}),$ (36)
@@ A.5 Baselines
FlowChef [38] Additionally, [26] is employed for FlowChef as well, using 200 NFE for inpainting and 50 NFE for all other tasks, a classifier-free guidance scale of 2, and a step size of 1 for all tasks.
@@ A.8 Additional Experimental Results
Table 3: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$ – Super-resolution ($\times 8$ and $\times 12$).
SR $\times 8$ SR $\times 12$
@@
A.9 FLAIR in Pixel Space
Table 5: Quantitative results with 50 NFE and $\sigma_{\nu}=0.5\%$ – In-painting and Super-resolution ($\times 8$).
Inpainting SR $\times 8$
[…]
DDNM 0.158 $\pm$ 0.042 26.9 0.732 $\pm$ 0.037 18.31 $\pm$ 2.94 0.199 $\pm$ 0.052 31.9 0.635 $\pm$ 0.079 23.59 $\pm$ 1.64
DPS 0.195 $\pm$ 0.064 30.2 0.689 $\pm$ 0.077 20.49 $\pm$ 2.81 0.172 $\pm$ 0.058 27.8 0.658 $\pm$ 0.088 24.59 $\pm$ 2.04
$M$ 0.161 $\pm$ 0.054 28.8 0.728 $\pm$ 0.062 20.59 $\pm$ 3.37 0.172 $\pm$ 0.051 29.1 0.669 $\pm$ 0.083 24.65 $\pm$ 1.97
[…]
FLAIR 0.097 $\pm$ 0.035 14.2 0.831 $\pm$ 0.031 21.87 $\pm$ 2.66 0.143 $\pm$ 0.039 22.9 0.712 $\pm$ 0.076 25.93 $\pm$ 1.96
[…]
Figure 2: Qualitative comparison. FLAIR in pixel space produces posterior samples of high perceptual quality while maintaining high data likelihood as well. Best viewed zoomed in.
[…]
A.10 Runtime Analysis
We compare the runtime and memory consumption of our method to the baselines. As our hard data consistency can strongly influence the runtime, we also provide measurements with the number of data term steps $\leq 5$ and additionally a fast version using a "tinyVAE" of SD3. To validate that the usage of the tinyVAE or less steps does not degrades the performance noticeably we also provide a metrics for x12 Super Resolution on 100 samples of FFHQ:
Table 6: Comparison of different methods in terms of runtime and memory usage. We validate the use of less data steps and a "tinyVAE" on $\times 12$ super resolution on 100 samples of the FFHQ dataset.
Method Runtime (s) $\downarrow$ Memory (MB) $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$
Resample 88.02 19009.2 0.461 25.31
FlowDPS 34.15 12228.6 0.404 27.74
RSD (no repulsion) 21.19 12400.0 0.462 28.11
FlowChef 15.23 12227.72 0.361 26.56
FLAIR (HDC, large VAE) 172.34 12389.4 0.259 27.42
FLAIR (HDC, tiny VAE) 40.77 5960.2 0.256 27.59
FLAIR (5 data term steps, tiny VAE) 22.46 5960.2 0.264 27.61
A.11 Ensembling Experiment
[…]
For ensembling we averaged 8 independent predictions of the corresponding methods. It can be seen that PSNR improves for all methods when ensembling. However, FLAIR shows the biggest gain, which means that our samples are indeed distributed around the mean and feature a higher variance compared to the baselines.
SR $\times 8$ SR $\times 12$
[…]
DIV2K
FlowDPS 22.53 0.4837 21.47 0.5524
FlowDPS (8x ensemble) 23.28 0.5157 22.06 0.5995
FlowChef 21.44 0.4898 20.20 0.5228
FlowChef (8x ensemble) 22.60 0.5502 21.60 0.5931
FLAIR 22.83 0.3627 21.05 0.4270
FLAIR (8x ensemble) 23.79 0.4244 22.27 0.4930
FFHQ
FlowDPS 29.02 0.3659 27.74 0.4036
FlowDPS (8x ensemble) 30.12 0.3267 28.65 0.3749
FlowChef 28.05 0.3303 26.56 0.3609
FlowChef (8x ensemble) 29.54 0.3267 28.15 0.3602
FLAIR 29.36 0.2028 27.42 0.2594
FLAIR (8x ensemble) 30.94 0.2457 29.00 0.2999
To highlight that our model can also be used to obtain good MMSE estimates, we also conducted ensemble predictions by running posterior sampling 8 times and averaging the result. The results show that ensembling increases PSNR values, but reduces LPIPS and confirms that our samples are indeed distributed around the posterior mean and that results very close to the posterior mean like the baseline methods are perceptually further away (LPIPS) from the ground truth compared to our samples.
A.12 Statistical Relevance
Our method is training-free, and the variance in reconstructed images is intentional, reflecting the stochasticity of our sampling process rather than instability. All methods are evaluated with identical random seeds to ensure fair comparison. We compute metrics over 1000 samples for FFHQ and 800 samples for DIV2K. Perceptual FID (pFID) is evaluated on $256\times 256$ patches, resulting in 9000 and 7200 samples, respectively.  Table 3 in Appendix A.8 reports means and standard deviations over multiple samples.
We also evaluated 100 FFHQ samples and 80 DIV2K samples, sampling three reconstructions per input for each method. We report the mean of each metric across all samples and the standard deviation of the means.
Table 8: Statistical evaluation on FFHQ for $\times$8 Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.370 $\pm$ 0.0012 70.7 $\pm$ 1.45 0.755 $\pm$ 0.0010 28.98 $\pm$ 0.008
RSD 0.4678 $\pm$ 0.0001 102.9 $\pm$ 0.05 0.7362 $\pm$ 0.0001 28.45 $\pm$ 0.001
FlowChef 0.3316 $\pm$ 0.0055 63.5 $\pm$ 1.45 0.7593 $\pm$ 0.0027 28.12 $\pm$ 0.072
Ours 0.2039 $\pm$ 0.0048 40.5 $\pm$ 0.94 0.7970 $\pm$ 0.0228 29.74 $\pm$ 0.668
Table 9: Statistical evaluation on FFHQ for $\times$12 Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.4073 $\pm$ 0.0002 77.6 $\pm$ 1.05 0.7391 $\pm$ 0.0006 27.71 $\pm$ 0.016
RSD 0.5039 $\pm$ 0.0002 119.0 $\pm$ 0.12 0.7217 $\pm$ 0.0001 27.08 $\pm$ 0.001
FlowChef 0.3626 $\pm$ 0.0050 81.1 $\pm$ 1.03 0.7283 $\pm$ 0.0027 26.62 $\pm$ 0.059
Ours 0.2593 $\pm$ 0.0023 45.8 $\pm$ 1.51 0.7582 $\pm$ 0.0252 27.81 $\pm$ 0.660
Table 10: Statistical evaluation on FFHQ for Motion Blur. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.4140 $\pm$ 0.0030 83.83 $\pm$ 1.00 0.7383 $\pm$ 0.0006 27.47 $\pm$ 0.05
RSD 0.4515 $\pm$ 0.0001 108.75 $\pm$ 0.08 0.7437 $\pm$ 0.0001 27.40 $\pm$ 0.00
FlowChef 0.4019 $\pm$ 0.0007 74.89 $\pm$ 0.69 0.7178 $\pm$ 0.0022 25.50 $\pm$ 0.04
Ours 0.2196 $\pm$ 0.0080 38.8 $\pm$ 2.25 0.7964 $\pm$ 0.0319 30.10 $\pm$ 0.96
Table 11: Statistical evaluation on FFHQ for Inpainting. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.3315 $\pm$ 0.0015 74.00 $\pm$ 0.58 0.7755 $\pm$ 0.0007 19.06 $\pm$ 0.11
RSD 0.4601 $\pm$ 0.0003 103.02 $\pm$ 0.03 0.7430 $\pm$ 0.0000 22.19 $\pm$ 0.01
FlowChef 0.3771 $\pm$ 0.0013 102.22 $\pm$ 0.26 0.7888 $\pm$ 0.0016 18.42 $\pm$ 0.25
Ours 0.1761 $\pm$ 0.0012 33.23 $\pm$ 2.02 0.8423 $\pm$ 0.0172 24.07 $\pm$ 0.80
Table 12: Statistical evaluation on DIV2K for $\times 8$ Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.5517 $\pm$ 0.0046 138.24 $\pm$ 1.67 0.5207 $\pm$ 0.0022 22.41 $\pm$ 0.02
RSD 0.7163 $\pm$ 0.0002 181.83 $\pm$ 0.15 0.4892 $\pm$ 0.0001 21.99 $\pm$ 0.00
FlowChef 0.5726 $\pm$ 0.0046 145.77 $\pm$ 3.31 0.4998 $\pm$ 0.0014 21.21 $\pm$ 0.04
Ours 0.3716 $\pm$ 0.0161 88.08 $\pm$ 1.43 0.5991 $\pm$ 0.0192 23.06 $\pm$ 0.41
Table 13: Statistical evaluation on DIV2K for $\times 12$ Super Resolution. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.6264 $\pm$ 0.0045 154.31 $\pm$ 0.43 0.4866 $\pm$ 0.0024 21.37 $\pm$ 0.02
RSD 0.7714 $\pm$ 0.0002 198.85 $\pm$ 0.12 0.4683 $\pm$ 0.0001 21.15 $\pm$ 0.00
FlowChef 0.6020 $\pm$ 0.0060 151.32 $\pm$ 2.45 0.4586 $\pm$ 0.0020 20.10 $\pm$ 0.03
Ours 0.4316 $\pm$ 0.0151 101.12 $\pm$ 4.51 0.5236 $\pm$ 0.0229 21.35 $\pm$ 0.51
Table 14: Statistical evaluation on DIV2K for Motion Deblur. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.6242 $\pm$ 0.0082 161.57 $\pm$ 3.90 0.4978 $\pm$ 0.0028 21.41 $\pm$ 0.04
RSD 0.8067 $\pm$ 0.0002 216.37 $\pm$ 0.36 0.4364 $\pm$ 0.0001 20.82 $\pm$ 0.00
FlowChef 0.6292 $\pm$ 0.0007 158.08 $\pm$ 3.03 0.4557 $\pm$ 0.0025 19.62 $\pm$ 0.05
Ours 0.3069 $\pm$ 0.0036 77.77 $\pm$ 1.89 0.6596 $\pm$ 0.0316 24.46 $\pm$ 0.71
Table 15: Statistical evaluation on DIV2K for Inpainting. We report mean $\pm$ standard deviation over 3 reconstructions per input.
Method LPIPS $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
FlowDPS 0.3738 $\pm$ 0.0032 106.58 $\pm$ 0.89 0.6579 $\pm$ 0.0009 21.06 $\pm$ 0.05
RSD 0.4667 $\pm$ 0.0003 136.82 $\pm$ 0.21 0.6650 $\pm$ 0.0001 23.07 $\pm$ 0.00
FlowChef 0.5111 $\pm$ 0.0019 128.97 $\pm$ 0.60 0.6355 $\pm$ 0.0006 20.51 $\pm$ 0.02
Ours 0.1729 $\pm$ 0.0014 51.41 $\pm$ 0.48 0.8122 $\pm$ 0.0124 24.06 $\pm$ 0.87
A.12.1 t-Test Analysis
We further performed paired t-tests on the LPIPS scores between FlowDPS and FLAIR. The null hypothesis states that the mean LPIPS scores are the same for both methods. In all settings, we reject the null hypothesis ($p<0.001$), confirming the statistical significance of our improvements see Table 16.
Table 16: Paired t-test $p$-values for LPIPS (FlowDPS vs. Ours). All comparisons are statistically significant.
Dataset Task p-value
[…]
Motion Deblur $5.30\times 10^{-4}$
Inpainting $1.21\times 10^{-4}$
[…]
Motion Deblur $4.29\times 10^{-5}$
Inpainting $3.00\times 10^{-5}$
A.13 Additional Qualitative Examples
@@
Figure 4: $\times 12$ super-resolution results on FFHQ. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
@@
Figure 7: $\times 12$ super-resolution results on DIV2k. Shown are observation, reference methods, FLAIR and ground truth. FLAIR produces sharp and results which still fulfill the data term, whereas the baselines tend to predict blurry images.
@@
Refer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSDRefer to captionRefer to captionRefer to captionRefer to captionFLAIRRefer to captionRefer to captionRefer to captionRefer to captionFlowDPSRefer to captionRefer to captionRefer to captionRefer to captionRSD